OpenStudy (anonymous):

Solving for X: X/36=X/40+12

OpenStudy (anonymous):

x\36=x/4+12 36/4=9 9+12 is 22 so x/22 x=1/22

OpenStudy (anonymous):

thx for ur help @ callie2240 Not what the answer key has, tho...and I'm off by a factor of 10.

OpenStudy (anonymous):

Oh, I'm sprry. Wish i could help you more